OpenStudy (anonymous):

Solve:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 3 }{ 3+2x^2 }\]

hartnn (hartnn):

do you know the standard result of \(\huge \int \dfrac{1}{x^2+a^2}dx=...?\) or if you don't want to use that you can do a trigonometric substitution to solve.

OpenStudy (anonymous):

yes i know it, but 3 isn't a square number :P

OpenStudy (anonymous):

and neither is 2. What substitution could i do??

OpenStudy (anonymous):

You can always force things :D \[\huge 3 = \sqrt3 ^2\]

hartnn (hartnn):

\(\large 3+2x^2 = 2(\dfrac{3}{2}+x^2) =2[ (\sqrt {\dfrac{3}{2}})^2+x^2] \) see what i did there ?

OpenStudy (anonymous):

oh, i see. Well, that'll be painful to do, but since when is math not :P

hartnn (hartnn):

math will not be painful anymore after you have enough practice ;)

OpenStudy (anonymous):

I love math, but squaring and rooting at the same time is not cool man.

OpenStudy (anonymous):

and @PeterPan, i'm loving the name ;)

OpenStudy (anonymous):

I love it too :)

OpenStudy (anonymous):

But sometimes, what seems to make things hard, actually makes things easier :D

OpenStudy (anonymous):

true dat.

OpenStudy (anonymous):

ok so now how do i solve \[\frac{ 1 }{ (x-3)^2 }\] I know it's a really stupid question, but i just can't remember how to do it :P

OpenStudy (anonymous):

FOIL? Or actually, more specifically, square of the binomial.

hartnn (hartnn):

you can substitute u= x-3

hartnn (hartnn):

du=dx and i think you know how to integrate 1/u^2 du

OpenStudy (anonymous):

omg i'm an idiot.